OpenStudy (anonymous):

Assume that all variables are nonzero Simplify 119x exponent24/7x exponent8

OpenStudy (anonymous):

I got 119/7x exponent16

OpenStudy (mathstudent55):

\( \dfrac{119x^24}{7x^8} \) Can you divide 119 by 7? what do you get? For the x part, use this rule. You got it correct. \( \dfrac{a^m}{a^n} = a^{m - n} \) Also, use the symbol ^ (shift 6) for exponent. x^8 means x to the 8 power.

OpenStudy (anonymous):

I tried to do that on the computer couldnt get it ,But I got my answer correct

OpenStudy (anonymous):

Assume that all variables are nonzero 28r^6s^7t^9/7r^3s^7t^6

OpenStudy (mathstudent55):

For the first problem, did you divide 119 by 7?

OpenStudy (mathstudent55):

\( \dfrac{28r^6s^7t^9}{7r^3s^7t^6} \) \( = \dfrac{28}{7} \times \dfrac{r^6}{r^3} \times \dfrac{s^7}{s^7} \times \dfrac{t^9}{t^6} \) Divide the numbers in the first fraction. For all the fractions with variables, apply the rule above where you subtract the exponents. What do you get?

OpenStudy (anonymous):

4=10/10

OpenStudy (mathstudent55):

What? 10/10 = 1, which is not equal to 4. What do you mean?

OpenStudy (anonymous):

Help me with this ...But back to the question it stated that assuming all the variable are zero

OpenStudy (mathstudent55):

All variable are NONzero.

OpenStudy (mathstudent55):

\(= \dfrac{28}{7} \times \dfrac{r^6}{r^3} \times \dfrac{s^7}{s^7} \times \dfrac{t^9}{t^6}\) \( = 4 \times r^3 \times s^0 \times t^3\) \(= 4r^3t^3\)

OpenStudy (anonymous):

y\[y^{?}/y ^{8}=y ^{9} Find the Missing Exponent \]

OpenStudy (mathstudent55):

\( \dfrac{y^?}{y^8} = y^9\) Apply the rule of division of powers: \( y^{? - 8} = y^9\) What number minus 8 equals 9?

OpenStudy (mathstudent55):

? - 8 = 9 Add 8 to both sides: ? = 17